A.        What is meant by the term: atomic structure?

B.       State Pauli's Exclusion principle

C.          Write the electronic configuration for the following:

         Neon [Ne - 10]

Argon [Ar - 18]

D.  Sodium chloride (NaCl) exhibits predominantly ionic bonding. The [Na - 11]+ and [Cl — 17]- ions have electron structures that are identical to which two inert gases?

                                                                                                                

E. Clearly show that the atomic packing factor of face-centred cubic structure is 0.74.


Calculate the radius of a barium atom, given that Ba has a BCC crystal structure, a density of 3.6.g/cm3 and an atomic weight of 137 g/mol.



1
Expert's answer
2022-01-18T16:18:01-0500

A. Atomic structure refers to the structure of an atom comprising a nucleus (centre) in which the protons (positively charged) and neutrons (neutral) are present. The negatively charged particles called electrons revolve around the centre of the nucleus.

B. Pauli's Exclusion Principle states that no two electrons in the same atom can have identical values for all four of their quantum numbers. In other words, (1) no more than two electrons can occupy the same orbital and (2) two electrons in the same orbital must have opposite spins.

C. Ne 1s22s22p6

Ar 1s22s22p63s23p6

D. Na+ 1s22s22p6 = Ne

Cl- 1s22s22p63s23p6 = Ar

E.


In BCC there are 2 atoms per unit cell, so

2/a3 = NA/Vmolar , where Vmolar = A/ρ; A is the atomic mass of Ba.

2/a3 = NA x ρ / A

a = (2A / NA x ρ)1/3 = (2 x 137 / 6.02 x 1023 x 3.6) = 5.02 x 10-8 cm